# ReDDiT: Rehashing Noise for Discrete Visual Generation

### v2

@@ Abstract
Discrete diffusion models are gaining traction in the visual generative area for their efficiency and compatibility. However, the pioneered attempts still fall behind the continuous counterparts, which we attribute to the noise (absorbing state) design and sampling heuristics. In this study, we propose the rehashing noise framework for discrete diffusion transformer (termed ReDDiT), to extend absorbing states and improve expressive capacity of discrete diffusion models. ReDDiT enriches the potential paths that latent variables traverse during training with randomized multi-index corruption. The derived rehash sampler, which reverses the randomized absorbing paths, guarantees the diversity and low discrepancy of the generation process. These reformulations lead to more consistent and competitive generation quality, mitigating the need for heavily tuned randomness. Experiments show that ReDDiT significantly outperforms the baseline (reducing gFID from 6.18 to 1.61) and is on par with the continuous counterparts with higher efficiency.
@@ 1 University of Chinese Academy of Sciences 2 China Mobile Research Institute
Project Page: https://github.com/martian422/ReDDiT
@@ 1 Introduction
Diffusion has been a highly competitive approach for generative workloads [8, 30, 21], offering strong bidirectional perception and well-structured mechanisms [48] for global control over content. In the continuous domain, diffusion transformers (DiTs) [28], which progressively refine image latents from Gaussian noise, have achieved impressive and scalable results. Despite the popularity of DiTs, the community shows a growing interest in discrete diffusion models in recent years [16, 40]. This trend is based on their practical advantages, e.g., compatibility with language models for the indexable codebooks and efficiency for the ability to generate multiple tokens at each inference. Early endeavors [4, 5, 12] mainly pursue efficiency through integrating visual tokenizers and BERT-style [mask] tokens [7]. Recent studies [2, 45] improved the generation quality, demonstrating great potential of visual generation in discrete domain.
@@
Despite the progress, discrete diffusion methods remain lagging behind their continuous counterparts. State-of-the-art approaches, e.g., masked visual token models (MVTMs) [45], is puzzeld by its mask design and predict-remask sampler, which restricts the model’s expressive capacity and makes the prediction performance sensitive to adaptions even with extensive training, Fig. 1(upper). Moreover, when paired with large-vocabulary codebooks from high-fidelity modern tokenizers, these models encounter challenges such as slower sampling speeds and numerical inaccuracy [49].
@@
Figure 1: Comparison of ReDDiT with the baseline discrete model (MVTM). MVTMs rely on score-based remasking strategies with Gumbel-max to sample from logits, which leads to lower token diversity and suboptimal token selection. In contrast, ReDDiT introduces a principled, low-discrepancy rehashing mechanism that leverages softmax-based probabilities, enabling diverse, high-quality sampling through a learned distribution. (Best viewed in color)
@@
To address these limitations, we propose two key hypotheses. First, while previous methods learn to recover plausible tokens from a monotonous [mask] canvas, this noise design may not be well-suited for discrete visual generation. In continuous diffusion, Gaussian noise is used to progressively degrade the input to learn a smooth distribution shift  [15, 22]. Discrete masking mimics this paradigm by collapsing all masked tokens into a single absorbing state, which however, lacks the variability of Gaussian noise, in terms of both vocabulary richness and latent diversity. Consequently, the discrete process offers a far coarser signal, which limits its ability to capture and represent diverse data distributions [33, 1]. Moreover, while continuous diffusion models introduce stochasticity at every inference step through noise injection, discrete unmasking is inherently binary: tokens are either masked or deterministically decoded. This rigid mechanism further constrains the flexibility of sample refinement during generation.
@@
Second, the predict-remask sampler of MVTMs introduces a form of handcrafted randomness, which is often implemented using Gumbel-max, to approximate sampling diversity. However, this sampler may compromise the probabilistic fidelity of the generation, and the need to carefully balance the token numbers decoded per step to mitigate accumulation errors leads to redundant sampling passes. As a result, the Gumbel-max has evolved to a heavily tuned time variant trick, leading to unstable performance, particularly when scaled to large-vocabulary codebooks. The above factors, rather than quantization alone, attribute to the performance gap between discrete and continous models.
@@
In this study, we propose a discrete diffusion model with an elaborate rehashing noise design, Fig. 1(lower). Our approach, termed ReDDiT, addresses the limitations of the uni-mask design by redefining absorbing states towards larger representational capacity, through enriching the potential paths that latent variables can traverse during diffusion. Specifically, we expand the masks to multiple indices along with the codebook and randomize them during data corruption. A rehash sampler is also derived with principled discrete diffusion theories to reverse the diffusion path for generation, guaranteeing the diversity and low discrepancy of the sampling process. We demonstrate that this rehashed noise helps learning a superior and flatter lower error bound while eliminating reliance to hyper-parameterized randomness during sampling.
@@
We further revisit the commonly used discrete diffusion objective (e.g., the masked cross-entropy loss) and update it with empirical modifications. By adopting an improved ELBO [31, 36] with representation alignment (RepA) [46] loss, we achieve better training efficiency and generation quality. Moreover, ReDDiT aligns with recent advances in discrete flow matching [11, 34], enabling token refreshment during sampling without training post-correction models [19].
2 Related Work
Diffusion Models.
Diffusion models [15, 37] have emerged as a powerful class of generative methods that learn data distributions by reversing a gradual noising process over time. These models are primarily designed for continuous domains such as images [8, 10, 28], defining a forward process that transforms data $x_{0}$ into noise $x_{1}$: $x_{t}\sim\mathcal{N}(\sqrt{\alpha_{t}}x_{0};(1-\alpha_{t})\mathbf{I})$ where $\alpha_{t}$ controls the noise schedule. The generative (reverse) process learns a denoising model $p_{\theta}(x_{s}\mid x_{t})$, often parameterized via a neural network $\theta$ to predict either noise or clean data.
Discrete Diffusion Models.
Discrete diffusion has been previously governed by masked visual token models (MVTMs) [4, 5, 12, 44, 45]. This model leverages a BERT-style [mask] token to corrupt the tokenized image sequence and trained the network with a simple cross-entropy loss on masked tokens, resulting in a score-based prediction. It generates tokens in a non-autoregressive fashion, by remasking the tokens with least scores at each inference as depicted in Alg. 1.
Recent studies unlocked the principled discrete diffusion model (DDM) [31, 36] and discrete flow-matching (DFM) [11, 34], which adapt the Markov chain theory, enabling generation over text [26, 23], molecules[34], and other discrete representations [1, 24]. Unlike MVTMs, the principled DDM and DFM mostly derive a time-weighted cross-entropy loss to supervise the training procedure and apply a gradual unmasking method based on probabilities.
@@ 3 Methodology
For self-containment, we first review the DDM theory in Sec. 3.1. We reformulate its diffusion dynamics and introduce rehashing noise for ReDDiT in Sec. 3.2. Connection and comparison of ours and other discrete diffusion models are discussed in Sec. 3.3.
@@ 3.1 Preliminary: Discrete Diffusion Model
DDM defines a forward process over discrete variables by gradually corrupting the image tokens into absorbing states (masks) through a continuous-time Markov process. Assume that the data consists of tokens from a finite vocabulary $\mathcal{V}$, $x\in\mathcal{V}^{L}$ represents a sequence of tokens ($e.g.$, an image tokenized into indices) with length $L$. We denote the clean data as $x_{t=0}$ ($x_{0}$ for short), and noise it gradually as $t\rightarrow 1$. DDM defines an absorbing token $\mathbf{m}\in\mathcal{V}$, such that once a token is noised to $\mathbf{m}$ it remains unchanged. At the terminal time $t=1$, $x_{t}$ fully transits to $\mathbf{m}^{L}$, which means $x_{1}^{i=1\sim L}=\mathbf{m}$.
Let $\alpha_{t}$ be the noise scheduler (a monotonically decreasing survival function that satisfies $\alpha_{0}=1,\alpha_{1}=0$ ). For $0\leq s<t\leq 1$, denote $\alpha_{t|s}^{\leftarrow}=\alpha_{t}/\alpha_{s}$, the forward corruption process is governed by a continuous-time transition kernel $q(x_{t}^{i}\mid x_{s}^{i})$ on the $i$-th element, as
$q(x_{t}^{i}\mid x_{s}^{i})=\begin{cases}1-\alpha_{t|s}^{\leftarrow},&\text{if }x_{t}^{i}=\mathbf{m},x_{s}^{i}\neq\mathbf{m},\\ \alpha_{t|s}^{\leftarrow},&\text{if }x_{t}^{i}=x_{s}^{i},x_{s}^{i}\neq\mathbf{m},\\ 1,&\text{if }x_{t}^{i}=x_{s}^{i},x_{s}^{i}=\mathbf{m},\\ 0,&\text{otherwise}.\end{cases}$ (1)
Given the transition kernel, the corrupted data distribution at time $t$ is determined  [31] as
$x_{t}\sim\alpha_{t}x_{0}+(1-\alpha_{t})\mathbf{m}^{L}$ (2)
The generative model learns the reverse process $p_{\theta}(x_{s}\mid x_{t})$, which denoises sample $x_{t}$ at arbitrary time $t\in(0,1]$ to a less noised state $x_{s}$ at time $s<t$. Denote $\delta(x_{t}^{i},\textbf{m})$ as the indicator function that only computes on masked tokens, and $\alpha_{t}^{\prime}=\frac{\mathrm{d}\alpha_{t}}{\mathrm{d}t}$, the learning objective is derived  [36, 31] as
$\mathcal{L}_{\text{DDM}}=-\mathbb{E}_{x_{0},\ x_{t}}\int_{t=0}^{t=1}[\frac{\alpha_{t}^{\prime}}{1-\alpha_{t}}\sum_{i=1}^{L}\delta(x_{t}^{i},\textbf{m})\log p_{\theta}(x_{0}^{i}\mid x_{t})]\mathrm{d}t\ .$ (3)
For a linear scheduler, Eq. 3 is simplified via variable substitution [36, 23] to an equivalent form:
$\mathcal{L}_{\text{DDM-linear}}=-\mathbb{E}_{t,\ x_{0},\ x_{t}}[\frac{1}{t}\sum_{i=1}^{L}\delta(x_{t}^{i},\textbf{m})\log p_{\theta}(x_{0}^{i}\mid x_{t})]\ .$ (4)
For conditional generation, class information $c$ (e.g., labels or text prompts) is introduced to the denoising model as additional input. Following classifier-free guidance [14], the model is trained with a random drop of labels, and the prediction is interpolated at inference:
@@
where $w\geq 0$ controls the guidance strength.
@@
3.2 Discrete Diffusion with Rehashing Noise
The ordinal structure inherent in discrete data provides a valuable inductive bias for designing transition kernels in diffusion dynamics. Prior studies [1, 3] show that assigning higher transition probabilities to neighboring pixel values—forming a discrete Gaussian-like noise—outperforms the single absorbing state approach on pixel-level datasets like CIFAR-10. However, when using visual tokenizers, the structure of discretized latents is learned rather than pre-defined, making such ordinal assumptions inapplicable. This insight motivates us to extend the conventional mask token into a set of indices, and reverse the diffusion path with noise rehashing. Such a procedure allows the model to optimize its embedding space during training, enhancing its ability to model flexible and data-driven noise structures. We visualize the learned distributions in Fig. 2 (right).
@@ 3.2 Discrete Diffusion with Rehashing Noise
Given $d$ categories, let $\mathbf{e}_{i}\in\mathbb{R}^{d}$ be its one-hot vector where the $i$-th value is $1$. We denote =$\mathcal{E}=\{\mathbf{e}_{i}\in\mathbb{R}^{d}\mid i=1,\ldots,d\}$ as the basis of a categorical distribution (known as $d$-simplex), and a basis for absorbing states with capacity $m$: =$\mathcal{M}=\{\mathbf{m}_{j}\in\mathbb{R}^{m}\mid j=1,\ldots,m\}$. The sum of $\mathcal{E}$ and $\mathcal{M}$ can be denoted as

### v3

@@ Abstract
In the visual generative area, discrete diffusion models are gaining traction for their efficiency and compatibility. However, pioneered attempts still fall behind their continuous counterparts, which we attribute to noise (absorbing state) design and sampling heuristics. In this study, we propose a rehashing noise approach for discrete diffusion transformer (termed ReDDiT), with the aim to extend absorbing states and improve expressive capacity of discrete diffusion models. ReDDiT enriches the potential paths that latent variables traverse during training with randomized multi-index corruption. The derived rehash sampler, which reverses the randomized absorbing paths, guarantees high diversity and low discrepancy of the generation process. These reformulations lead to more consistent and competitive generation quality, mitigating the need for heavily tuned randomness. Experiments show that ReDDiT significantly outperforms the baseline model (reducing gFID from 6.18 to 1.61) and is on par with the continuous counterparts. The code and models will be publicly available.
@@ 1 Introduction
Diffusion has been a competitive approach for generative workloads Dhariwal & Nichol (2021); Rombach et al. (2022b); Li et al. (2024), offering strong bidirectional perception and well-structured mechanisms Zhang et al. (2023) for global control over content. Within the continuous domain, diffusion transformers (DiTs) Peebles & Xie (2023), which progressively refine image latents from Gaussian noise, have achieved impressive and scalable results. Recently, the community shows a growing interest in discrete diffusion models Hu & Ommer (2024); Swerdlow et al. (2025), which is based on their practical advantages, e.g., compatibility with language models for the indexable codebook, and efficiency for predicting multiple tokens at each inference. Early endeavors Chang et al. (2022; 2023); Gu et al. (2022) pursue efficiency through integrating visual tokenizers and BERT-style [mask] tokens Devlin et al. (2019). Recent studies Bai et al. (2025); Yang et al. (2025) improved the generation quality, demonstrating great potential of discrete diffusion.
@@
Despite the progress, the performance of discrete diffusion methods remains lagging behind their continuous counterparts. Representative approaches, e.g., masked visual token models (MVTMs) Chang et al. (2022); Yu et al. (2023), are puzzled by the mask design and confidence-based re-mask sampler, which restricts model’s expressive capacity and makes prediction sensitive to adaptions given extensive training, Fig. 1(upper). Moreover, when paired with large-vocabulary codebooks from high-fidelity modern tokenizers, they encounter challenges such as slower sampling speeds and numerical inaccuracy Zheng et al. (2024).
@@
Figure 1: Comparison the baseline discrete model (MVTM) with ReDDiT. MVTMs rely on score-based remasking strategies with Gumbel-max to sample from logits, which leads to lower token diversity and suboptimal token selection. In contrast, ReDDiT introduces a systematic, low-discrepancy rehashing mechanism that leverages softmax-based probabilities, enabling diverse, high-quality sampling through a learned distribution. (This figure is best viewed in color)
@@
To address these limitations, we first propose two hypotheses. First, while discrete methods learn to recover plausible tokens from a monotonous [mask] canvas, the used noise design may not be well-suited for discrete visual generation. In continuous diffusion, Gaussian noise is used to progressively degrade the input to learn a smooth distribution shift  Ho et al. (2020); Lu et al. (2022). Discrete masking mimics this paradigm by collapsing all masked tokens to a single absorbing state, which, however, lacks the variability of Gaussian noise, in terms of both vocabulary richness and latent diversity. Consequently, the discrete process offers a far coarser signal, which limits its ability to represent diverse data distributions Santos et al. (2023); Austin et al. (2021). Moreover, while continuous diffusion models introduce stochasticity at every inference step through noise injection, discrete unmasking is inherently binary: tokens are either masked or deterministically decoded, Fig. 1(upper). This rigid mechanism constrains the flexibility of sample refinement during generation.
@@
Second, the confidence-based re-mask sampler of MVTMs introduces a form of handcrafted randomness, which is implemented through Gumbel-max, to approximate sampling diversity. Unfortunately, this sampler compromises the probabilistic fidelity of generation, and the need to carefully balance token numbers decoded per step (for mitigating accumulation errors) leads to redundant sampling passes. As a result, Gumbel-max has evolved to a heavily tuned time variant trick with unstable performance, particularly when scaled to large-vocabulary codebooks. The above factors, rather than quantization alone, induce the performance gap between discrete and continuous models.
@@
In this study, we propose a discrete diffusion model with an elaborate rehashing noise design, Fig. 1(lower). Our approach, termed ReDDiT, addresses the limitations of the uni-mask design by redefining absorbing states towards larger representational capacity, through enriching the potential paths that latent variables can traverse during diffusion. Specifically, we expand the masks to multiple indices along with the codebook and randomize them during data corruption. A rehash sampler is also derived with principled discrete diffusion theories to reverse the diffusion path for generation, guaranteeing high diversity and low discrepancy of the sampling process. We demonstrate that this rehashed noise facilitates learning a superior and regularized expressiveness, while eliminating reliance to hyper-parameterized randomness during sampling.
@@
We further revisit the commonly used discrete diffusion objective and update it with empirical modifications. By adopting an improved ELBO Sahoo et al. (2024); Shi et al. (2024) with representation alignment (RepA) Yu et al. (2025) loss, we optimize the training efficiency and substantially improve the generation quality of discrete generative models. Moreover, ReDDiT aligns with recent advances in discrete flow matching Gat et al. (2024); Shaul et al. (2024), enabling token refreshment during sampling without training post-correction models Lezama et al. (2022).
@@ 2 Methodology
For self-containment, we first review the DDM theory in Sec. 2.1. We then reformulate its diffusion dynamics and introduce rehashing noise for ReDDiT in Sec. 2.2. We finally discuss connection and comparison with other discrete diffusion models in Sec. 2.3.
@@ 2.1 Preliminary: Discrete Diffusion Model
DDM defines a forward process over discrete variables by gradually corrupting the image tokens to absorbing states (masks) through a continuous-time Markov process. Assume that the data consists of tokens from a finite vocabulary $\mathcal{V}$. $x\in\mathcal{V}^{L}$ is a sequence of tokens ($e.g.$, an image tokenized into indices) with length $L$. We denote the clean data as $x_{t=0}$ ($x_{0}$ for short), and noise it gradually as $t\rightarrow 1$. DDM defines an absorbing token $\mathbf{m}\in\mathcal{V}$, such that once a token is noised to $\mathbf{m}$ it remains unchanged. At the terminal time $t=1$, $x_{t}$ fully transits to $\mathbf{m}^{L}$, which means $x_{1}^{i=1\sim L}=\mathbf{m}$.
Let $\alpha_{t}$ be the noise scheduler (a monotonically decreasing survival function that satisfies $\alpha_{0}=1,\alpha_{1}=0$ ). For $0\leq s<t\leq 1$, the forward corruption process is governed by a continuous-time transition kernel $q(x_{t}^{i}|x_{s}^{i})$ at the $i$-th element, as
$q(x_{t}^{i}|x_{s}^{i})=\begin{cases}1-\alpha_{t|s},&\text{if }x_{t}^{i}=\mathbf{m},x_{s}^{i}\neq\mathbf{m}\\ \alpha_{t|s},&\text{if }x_{t}^{i}=x_{s}^{i},x_{s}^{i}\neq\mathbf{m}\\ 1,&\text{if }x_{t}^{i}=x_{s}^{i},x_{s}^{i}=\mathbf{m}\\ 0,&\text{otherwise}\end{cases},\quad\alpha_{t|s}=\frac{\alpha_{t}}{\alpha_{s}}.$ (1)
Denoting $q$ as the transition kernel and $\text{Cat}(\cdot;\pi)$ the categorical distribution determined by probability $\pi$, the corrupted data distribution at time $t$ is written as
$x_{t}\sim q(x_{t}|x_{0}),q(x_{t}|x_{0})=\text{Cat}(x_{t};\alpha_{t}x_{0}+(1-\alpha_{t})\mathbf{m}^{L}).$ (2)
The generative model learns the reverse process $p_{\theta}(x_{s}|x_{t})$, which denoises sample $x_{t}$ at arbitrary time $t\in(0,1]$ to a less noised state $x_{s}$ at time $s<t$. Denoting $\delta(x_{t}^{i},\textbf{m})$ as the indicator function that only computes on masked tokens, and $\alpha_{t}^{\prime}=\frac{\mathrm{d}\alpha_{t}}{\mathrm{d}t}$, the learning objective is derived  Shi et al. (2024) as
$\mathcal{L}_{\text{DDM}}=-\mathbb{E}_{x_{0},\ x_{t}}\int_{t=0}^{t=1}[\frac{\alpha_{t}^{\prime}}{1-\alpha_{t}}\sum_{i=1}^{L}\delta(x_{t}^{i},\textbf{m})\log p_{\theta}(x_{0}^{i}|x_{t})]\mathrm{d}t\ .$ (3)
For a linear scheduler, Eq. 3 is simplified via variable substitution Sahoo et al. (2024) to an equivalent form, as
$\mathcal{L}_{\text{DDM-linear}}=-\mathbb{E}_{t,\ x_{0},\ x_{t}}[\frac{1}{t}\sum_{i=1}^{L}\delta(x_{t}^{i},\textbf{m})\log p_{\theta}(x_{0}^{i}|x_{t})]\ .$ (4)
For conditional generation, class information $c$ (e.g., labels or text prompts) is introduced to the denoising model as additional input. Following classifier-free guidance Ho & Salimans (2022), the model is trained with a random drop of labels, and the prediction is interpolated at inference, as
@@
where $\varnothing$ is the dropped label and $w\geq 0$ controls the guidance strength.
2.2 Discrete Diffusion with Rehashing Noise
The ordinal structure inherent in discrete data provides a valuable inductive bias for designing transition kernels in diffusion dynamics. Prior studies Austin et al. (2021); Campbell et al. (2022) show that assigning higher transition probabilities to neighboring pixel values—forming a discrete Gaussian-like noise—outperforms the single absorbing state approach on pixel-level datasets like CIFAR-10. However, when using visual tokenizers, the structure of discretized latents is learned rather than pre-defined, making such ordinal assumptions inapplicable. This insight motivates us to extend conventional mask tokens to a set of indices, and reverse the diffusion path with noise rehashing. This design allows the model to optimize its embedding space during training, enhancing its ability to model flexible and data-driven noise structures. We visualize the learned distributions in Fig. 2 (right).
@@ 2.2 Discrete Diffusion with Rehashing Noise
Given $d$ categories, let $\mathbf{e}_{i}\in\mathbb{R}^{d}$ be its one-hot vector where the $i$-th value is $1$. We denote $\mathcal{E}=\{\mathbf{e}_{i}\in\mathbb{R}^{d}\mid i=1,\ldots,d\}$ as the basis of a categorical distribution (known as $d$-simplex), and a basis for absorbing states with capacity $m$: $\mathcal{M}=\{\mathbf{m}_{j}\in\mathbb{R}^{m}\mid j=1,\ldots,m\}$. The sum of $\mathcal{E}$ and $\mathcal{M}$ can be denoted as